OpenStudy (anonymous):

A wire bent into the shape of a square encloses an area of 25 cm square. Then the same wire is cut and bent into two identical circles. What is the radius of one of the circles? Round to the nearest hundreth.

OpenStudy (britnee_paige99):

Area =π(r^2) so you know that their are two circles and that they equal 25 cm^2 therefore you can divide by 2 and you know one circle has an area of 12.5. divide by pie using the equation and then you have 12.5/3.14=r^2. then you can square root both sides. (12.5/3.14)^(1/2)=r. to the 1/2 power is the same as the square root. Then you get the answer of 1.999. which is about 2 cm radius.

OpenStudy (anonymous):

Thank you!